OpenStudy (anonymous):

Olivia is 18 and would like to buy a house when she is 30. What is the discount factor for today’s prices if the housing values increase 8% per year?

hero (hero):

More information is most likely required in order to help you.

OpenStudy (anonymous):

Thats what i thought but thats all the question says :/

hero (hero):

You should go back and re-read the chapter of the book and find out the relationship between "discount factor" and "housing value increases".

OpenStudy (anonymous):

Why isn't it just: \[1/(1+r)^{t}\] or \[1/(1+.08)^{30-18}\] \[=1/(1.08)^{12}\] = .39711

OpenStudy (anonymous):

WHAT FORMULA IS THAT? thats right.

OpenStudy (anonymous):

It's been awhile since I did this. But I just know it as the formula for a discount value using simple interest. r = interest rate t = time (in this case 12 years until she buys a house)

OpenStudy (anonymous):

should have wrote above...formula for a discount factor using simple interest

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Discounting

OpenStudy (anonymous):

that formula assumes annual compounding of interest... Not sure what you are up to, but if you assumed continuous compounding of interest, the formula would be: \[e^{-rt}\]